OpenStudy (anonymous):

solve ln3-ln(x+5)-lnx=0

OpenStudy (turingtest):

\[ \ln 3-\ln(x+5)-\ln x=0\]try to first combine this into one log like we did earlier, what do you get?

OpenStudy (anonymous):

ln 3 ___ x+5 / x ??

OpenStudy (turingtest):

almost, but realize that\[\ln({3\over x+5})-\ln x=\ln({3\over x^2+5x})=0\]you just need to be more careful with the algebra. After you get that, raise the log's base to the power of both sides. Can you do that?

OpenStudy (turingtest):

If this step is giving you trouble, look at just the algebra:\[{{3\over x+5}\over x}={3\over x+5}*{1\over x}={3\over x(x+5)}={3\over x^2+5x}\]

OpenStudy (anonymous):

ok seeing it like that is much better, so from there i can just factor out?

OpenStudy (turingtest):

form here we have\[\ln({3\over x^2+5x})=0\]do you have any idea how to proceed from here? (hint: it's not factoring)

OpenStudy (anonymous):

quadratic formula?

OpenStudy (anonymous):

TuringTest: I admire you, taking your time to explain! Nice going!

OpenStudy (turingtest):

Not yet, we need to get rid of the log first. So we raise both sides to the power of the base, because\[a^{\log_a x}=x\] so, first question: what is the base if this log? (i.e. what is "a" for us?) @alifetouni thanks :)

OpenStudy (turingtest):

Sorry, typo, we raise the base of the log to the power of both sides :/ Terminology can be a little tricky with logs... So the question I am asking is what is the base of the logarithm \[\ln x\]?

OpenStudy (anonymous):

is it e

OpenStudy (turingtest):

yes! now what is\[e^{\ln({3\over x^2+5x})}=e^0\]?

OpenStudy (anonymous):

well the ln and e = 1 correct? so then you just have 3/x^2+5x=1 because anything raised to 0 is 1 right?

OpenStudy (turingtest):

perfect :) so now we have\[{3\over x^2+5x}=1\]can you solve this quadratic equation for x?

OpenStudy (anonymous):

umm im not to sure haha let me try

OpenStudy (turingtest):

Go for it! Hint: Try to get it looking like\[ax^2+bx+c=0\]then you can either factor it or use the quadratic formula.

OpenStudy (anonymous):

so i got x^2+5x-3=0

OpenStudy (turingtest):

good! This is not easily factorable, so can you apply either the quadratic formula or complete the square?

OpenStudy (anonymous):

-5+or-√37 ___________ 2

OpenStudy (turingtest):

That's what I got =) Good job, hope that helped. I have to go for a bit. ttyl...

OpenStudy (anonymous):

ok thanks a bunch!